# Extreme Compression of Large Language Models via Additive Quantization

### v2

@@ Appendix G Estimating model size
Appendix H End-to-End Inference Speed
@@
Original (float16) 41.51 26.76 5.66
AQLM (Table 1) 32.22 25.04 5.65
AQLM (2×2$\times$2 ×8-bit) 32.64 25.35 6.03
@@
For quantized Llama-2 models, the same setup as in Section 4.3, we measure the time it takes to generate 128 tokens from scratch with batch size 1, and report the average number of generated tokens per second on a single 24GB RTX 3090 GPU, as well as Intel i9 CPU, in Table 14.
Appendix I Codebook and codes distribution
The proposed AQLM quantization method allows for large freedom in the choice of quantization lattice and ability to represent different weight distribution. To understand how do the learned codes and codebooks look like, we visualize the distribution of codes (how frequently given codebook vector is chosen) and the learned codebooks. Below on Figure 7 we provide a histogram of leaned codes and two leading principal components for a specific layer. One can observe that codes are assigned uniformly across codebook vectors and codebook vectors are concentrated in some ball. This pattern is pertinent to all linear projections inside transformer blocks.
Figure 7: Visualization of learned codes and codebooks in layers.31.mlp.down_proj linear projection.

### v3

@@ Appendix H Estimating model size
Appendix I End-to-End Inference Speed
@@
Original (float16) 54.2 29.5 5.8
AQLM (1×1$\times$1 ×16-bit) 65.3 34.1 6.7
AQLM (2×2$\times$2 ×8-bit) 114.1 68.1 14.3
@@
For quantized Llama 2 models, setup described in Section 4.4, we measure the time it takes to generate 128 tokens from scratch, performed on compiled computational graphs, with batch size 1, and report the average number of generated tokens per second on a single 24GB RTX 3090 GPU, as well as Intel i9 CPU, in Table 14. Perplexity on WikiText-2 on these configurations presented at the Table 9
Appendix J Codebook and codes distribution
The proposed AQLM quantization method allows for large freedom in the choice of quantization lattice and ability to represent different weight distribution. To understand how do the learned codes and codebooks look like, we visualize the distribution of codes (how frequently given codebook vector is chosen) and the learned codebooks. Below on Figure 7 we provide a cumulative probability plot of leaned codes and two leading principal codebook components for a specific layer. One can observe that codes distribution is close to uniform. Its entropy equals 15.91 bits per code, which is close to the maximum possible entropy of 16 bits (for a 16-bit codebook) for the uniform distribution. Codebook vectors are concentrated in some ball. This pattern is pertinent to all linear projections inside transformer blocks.
Figure 7: Visualization of learned codes and codebooks in layers.5.self_attn.q_proj linear projection.
@@
Appendix K Evaluation on MMLU and GSM8k
While measurement of perplexity on WikiText-2 and C4 together with zero-shot accuracy on subset of simple 0-shot tasks from LM Eval Harness (Gao et al., 2021) is an established benchmark for evaluation of performance of compressed models, it may be not exhaustive enough for many real-world cases. While the complete and exhaustive evaluation of LLM abilities is still an open question, we evaluate our AQLM models and QuIP# on MMLU (Hendrycks et al., 2020) benchmark that involves problems from 57 different domains, such as humanities, social sciences, physics, e.t.c, and GSM8k (Cobbe et al., 2021) to assess the performance of quantized models on more complex and challenging tasks, requiring reasoning to get the correct answer. Below we consider AQLM and QuIP# after end-to-end finetuning, i.e. the best performing quantized models. We observed that relative decrease on performance on these tasks is higher compared to the standard evaluation. Fine-tuned AQML and QuIP# yield very similar performance on these benchmarks.
Table 15: Evaluation of quantized Llama 2 models for 2-2.1 bits per parameter on MMLU and GSM8k.
corresponds to end-to-end finetuning
Size Method Avg bits MMLU (5-shot) GSM8k (8-shot)
7B 16 45.9 14.6
QuIP# 2.02 36.8 6.2
AQLM 2.02 38.5 5.3
13B 16 55.2 24.3
QuIP# 2.01 50.0 14.0
AQLM 1.97 48.8 13.8
70B 16 68.8 56.3
QuIP# 2.01 65.3 46.4
AQLM 2.07 65.3 47.9
Appendix L Block-wise tuning for scalar quantization
The block-wise procedure introduced in our work is quite general and can be applied to scalar quantization as well. Specifically, operations with quantized weights are differentiable with respect to quantization scales kept in original precision. Therefore, scales can be tuned in the same way as AQLM codebooks. We observed that tuning significantly improves the quality of GPTQ at low bit widths. However, the resulting quality is still far below AQLM at similar bit-widths.
Table 16: Evaluation of AQLM and GPTQ quantization after block tuning for Llama 2 models with 2-2.1 bits per parameter.
Size Method Avg bits Wiki2$\downarrow$ C4$\downarrow$
7B 16 5.12 6.63
GPTQ 2.14 16.77 17.53
AQLM 2.02 6.64 8.56